IRTC removal

After shimming is completed the IRTC is decoupled from the AGW unit and stored away

Alignment verification

We double check the small bar mirror, to verify that it has not moved during alignment. We use the LT to compare the LUCI-facing IRTC flange with the small bar mirror, and the result is 0.0045 degrees = 16 arcsec. This is higher than what we had two days ago (0.0016 degrees) but still in spec.

Pennina work

To align the pennina we need the neon laser. The laser is found in the AO4 cabinet (while the spare list said AO3), but the power supply wants a 220V input. After a long search, we find the transformer is in use by ARGOS, and actually they have replaced it with a new model.

The tilt on the vertical axis is blocked by a teflon washer put between the pennina and the flowerpot case. We remove the teflon as a test, and we are able to move the pennina by about 1 degree (that would be 1/4 of the beam size). The teflon washer is machined to reduce its thickness by about half,

-- AlfioPuglisi - 07 Feb 2016
Topic revision: r1 - 07 Feb 2016, AlfioPuglisi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback